OpenStudy (anonymous):

csc(x)cot(x)(1-cos^2(x))=___x Can anyone help me?

OpenStudy (anonymous):

I'm not sure how to use identities to solve this, but that seems to be what it is asking me to do. If anyone can show me step by step how to solve this it would be great.

OpenStudy (anonymous):

I think I have gotten to \[1/\sin(x)((1/\cos(x))-\cos(x))\]

OpenStudy (loser66):